Meldrick is playing foot jump. The ball in his foot jump is in uniform


circular motion and makes 10 revolutions in 4.0 seconds.


a) What is its period?



b) If the length of the plastic cord that holds the ball is 0.8 meter, what is its tangential velocity?



c) If the ball has a mass of 3 grams, how much force is acting the ball to keep it in uniform circular motion?



d) What is the centripetal acceleration of the ball?

1
Expert's answer
2022-05-28T08:22:19-0400

Given:

"N=10\:\rm rev"

"t=4.0\:\rm s"

"R=0.8\:\rm m"

"m=0.003\:\rm kg"

(a)

"T=\frac{t}{N}=\frac{4.0}{10}=0.40\rm \:s"

(b)

"v=\frac{2\pi R}{T}=\frac{2\pi*0.8}{0.40}=13\:\rm m/s"

(c)

"F=ma=\frac{mv^2}{R}=\frac{0.003*13^2}{0.8}=0.59\:\rm N"

(d)

"a=\frac{v^2}{R}=\frac{13^2}{0.8}=200\:\rm m/s^2"